OpenStudy (anonymous):

A prize fund is set up with a single investment of $5000 to provide an annual prize of $500. The fund is invested to earn interest at a rate of 7% compounded annually. If the first prize is awarded one year after the initial investment, find the number of years for which the prize can be awarded before the fund falls below $500.

OpenStudy (nurali):

PV of n year annuity 5000=500*(1-1/1.07^n)/0.07 5000=7143*(1-1/1.07^n) 0.70=(1-1/1.07^n) 1/1.07^n=0.3 3.33=1.07^n taking 'ln' both side Ln(3.33)=Ln(1.07^n)=(n)Ln(1.07) n=Ln(3.33)/Ln(1.07) n=1.20/0.068 n=17.7 17 years; not enough money in year 18

OpenStudy (nurali):

Welcome to Openstudy.